OpenStudy (anonymous):

71/9-48/9= I dont understand this?

OpenStudy (anonymous):

If you have 71 oranges and you subtract 48 oranges, how many oranges do you have left?

OpenStudy (anonymous):

This is quite an algebraic manipulation. It is processed as shown If you have the following situation where the denominators are the same, you can just keep the same denominator and deal with the numerator as shown \[ \frac{a}{b} +\frac{c}{b} =\frac{a+c}{b} \]

OpenStudy (anonymous):

first you divide 71 to 9 and it is 8/1 and then you divide 41 devided by 9 and it is 16/1 and the answer is 2/1 that equals to 2

OpenStudy (anonymous):

\[(71/9)-(48/9)=71-48/9 = 23/9\]